OpenStudy (anonymous):

Write an equation for the circle with center (-4, 4) and containing the point (-1, 16). Use the ^ key for the exponents. Write your answer as the example: (x+2)^2+(y-9)^2=100

OpenStudy (anonymous):

is the point (-1 , 16) on the circle or within it? i cant tell from your question

OpenStudy (anonymous):

within it, i think

OpenStudy (anonymous):

(x+4)^2 + (y-4)^2 = r^2 where r^2 > 153

OpenStudy (anonymous):

(x+4)^2 +(y-4)^2=153

OpenStudy (anonymous):

it said its wrong..

OpenStudy (anonymous):

sorry, i should have explained more, 153 is the radius squared, what we do is find the distance between the centre and the point (-1 , 16), and that is the radius

OpenStudy (anonymous):

what does it say is the right answer?

OpenStudy (anonymous):

it doesnt lol it's a fill in the blank and i cant find it anywhere on the internet

OpenStudy (anonymous):

ah is this an online homework? im pretty sure that is the equation for the circle centre (-4, 4) and passing through point (-1 ,16)

OpenStudy (anonymous):

yea it is online and i guess so lol